OpenStudy (anonymous):

g(x) = √x2-3, find domain and range

OpenStudy (zzr0ck3r):

is the sqrt include all?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

is that x^2 or 2x?

OpenStudy (zzr0ck3r):

x^2-3 >= 0 x^2 >= 3 x>= sqrt 3 for domain^

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

then?

OpenStudy (zzr0ck3r):

well and x <= -sqrt(3)

OpenStudy (anonymous):

why negative?

OpenStudy (zzr0ck3r):

range is 0 to infinity

OpenStudy (zzr0ck3r):

because -sqrt(3) or less squared is more than 3 thus never negative under the radicant

OpenStudy (zzr0ck3r):

http://www.wolframalpha.com/input/?i=graph+sqrt%28x%5E2-3%29 ignore the red (complex) line

OpenStudy (anonymous):

okay, thanks :)

OpenStudy (zzr0ck3r):

the negative part comes from the line where x^2 = 3, so (+ or -) x>= sqrt3 then we have two equations x>= sqrt(3) and then we change the sign when we multiply the RHS to get the negative that we must retain so x <= -sqrt(3)

OpenStudy (anonymous):

Domain: What x is allowed to be Range: What y is allowed to be (Here y is renamed as g(x) ) I will assume you are working with the real numbers You have \[g(x)=\sqrt(x^2-3)\] But the inside of your radical must be greater than or equal to 0. In other words the inside of the radical is 0 or positive. In other words, you do not want to run into the problem that you will take the square root of a negative number (***because you are working with the real numbers) So for the domain you have to solve: \[x^2-3\ge0\] this is the inside of the radical, which says it's 0 or positive \[x^2\ge3\] Added 3 to both sides \[x \ge \sqrt3\] Took the square root of both sides So the domain is exactly this last inequality. That is what x is allowed to be. For range you simply start plugging in x's and see what you get for g(x) [or y]. when x=3 you get g(x)=0 And when you plug in bigger x's you find out g(x) simply increases, or gets bigger. So your range is \[g(x) \ge 0\]

OpenStudy (zzr0ck3r):

the last part is not right the domain will include anything smaller and -sqrt(3)

OpenStudy (anonymous):

You are correct zzr0ck3r. I forgot to include in the domain something that is very important. So from the equation \[x^2 \ge 3\] We will get \[x \ge \sqrt3\] or \[x \le -\sqrt3\] @AmiraZainal to understand this a bit easier let's try using nicer numbers: if we had ended up with \[x^2 \ge 9\] Your answer for x is which number when I multiply it by itself gives me at least 9? And you will realize that when x=3 you get at least 9 since 3*3=9, so you get \[x \ge 3\] But if you try -3 * -3 you will also get at least 9 So your answer also includes \[x \le -3\]